OpenStudy (anonymous):

WHICH BINOMIAL IS A FACTOR OF THE POLYNOMIAL? X^2+5-36

OpenStudy (anonymous):

mathway.com

OpenStudy (dan815):

binomial is a term with 2 terms

OpenStudy (dan815):

X^2+5-36=x^2-31

OpenStudy (dan815):

do u mean x^2+5-36

OpenStudy (dan815):

x^2+5x*-36

OpenStudy (anonymous):

YEAH

OpenStudy (dan815):

okay then u are solving this problem

OpenStudy (dan815):

a*b=-36 a+b=5

OpenStudy (dan815):

can you solve that?

OpenStudy (anonymous):

NOPE THATS WHY I CAME HERE.

OpenStudy (dan815):

i mean this a*b=-36 a+b=5

OpenStudy (anonymous):

I'll Help

OpenStudy (dan815):

can u do that problem

OpenStudy (dan815):

9*-4=-36 9+-4=5

OpenStudy (dan815):

make sense?

OpenStudy (dan815):

(x+9)(x-4)=x^2+9x-4x-36

OpenStudy (anonymous):

YEAH

OpenStudy (dan815):

binomial is an expression with 2 terms

OpenStudy (dan815):

x+9 is a binomial expression (x-4) is also a binomial expression